### v1

# Prompt Engineer: Analyzing Skill Requirements in the AI Job Market

Abstract
The rise of large language models (LLMs) has created a new job role: the Prompt Engineer. Despite growing interest in this position, we still do not fully understand what skills this new job role requires or how common these jobs are. We analyzed 20,662 job postings on LinkedIn, including 72 prompt engineer positions, to learn more about this emerging role. We found that prompt engineering is still rare (less than 0.5% of sampled job postings) but has a unique skill profile. Prompt engineers need AI knowledge (22.8%), prompt design skills (18.7%), good communication (21.9%), and creative problem-solving (15.8%) skills. These requirements significantly differ from those of established roles, such as data scientists and machine learning engineers, showing that prompt engineering is becoming its own profession. Our findings help job seekers, employers, and educational institutions in better understanding the emerging field of prompt engineering.
Keywords: Prompt engineering, large language models, LLM, AI workforce, skill analysis, job market analysis, text mining, natural language processing, human-AI interaction
1 Introduction
The rise of large language models (LLMs) has been transforming various industries (Chkirbene et al., 2024). AI enables advancements in automation, content generation, and AI-driven decision-making. However, the effectiveness of AI models heavily depends on how prompts are structured (Oppenlaender, 2023), leading to the emergence of prompt engineering as a specialized field. Companies are increasingly seeking to recruit prompt engineers in an effort to optimize and extract the full potential of LLMs. Despite the growing demand for this role, there remains no standardized definition of a prompt engineer, resulting in uncertainties regarding the scope, required skills, and industry relevance of this emerging job role.
Due to the timely relevance of LLM applications in industry, it is important to develop an understanding of the required knowledge and skills for this emerging job type. However, because of the novelty of AI technology, individuals who are interested in pursuing the role of the prompt engineer may face challenges understanding the qualifications required. Therefore, addressing this gap is beneficial to both job seekers and practitioners in industry as well as employers who want to identify suitable candidates. However, how can we systematically extract and analyze the skills, competencies, and industry relevance of this job role? In this paper, we employ a text mining approach to analyze job advertisements published on LinkedIn, aiming to uncover the key knowledge areas and skills required for prompt engineers. Understanding these trends provides insight into the broader adoption of LLMs and the pace of different industries in adapting to this evolving technology.
This research systematically analyzes job postings to develop a data-driven understanding of the prompt engineer profession. We specifically seek to answer the following research questions:
  • RQ1:
    How prevalent are prompt engineer positions compared to more established job roles?
  • RQ2:
    What soft skills are most valued for prompt engineers, and how do these differ from other job roles?
  • RQ3:
    What technical skills define the prompt engineer role, and how do these hard skills create a distinct professional profile?
By conducting a systematic quantitative analysis of job postings, this study aims to provide a snapshot of the current state of the prompt engineer profession, and valuable insights into the evolving landscape of prompt engineering and its implications for both job seekers and employers.
The remainder of this paper is structured as follows. We review related literature on prompt engineering and the extraction of skills in Section 2. We then describe our approach to collecting and extracting information from job descriptions in Section 3. We then present our findings in Section 4, contrasting the skills of prompt engineers with fixe other related jobs. We discuss our findings in Section 5 and conclude in Section 6.
2 Related Work
2.1 The Emergence of the Prompt Engineer
Artificial intelligence (AI) is changing the way workers perform their jobs and the skills they require (OECD, 2024; Brasse et al., 2024). Generative artificial intelligence has a strong impact on the future of work (Cazzaniga, 2024; Carroll, 2025) and higher education (Fan and Irawan, 2025). While concerns and anxiety over automation (Dinnen, 2022) are growing, AI also creates new professional opportunities (Elrawy and Wagdy, 2025; Kutela et al., 2023). There is a demand for a new digital workforce with a novel set of AI literacy (Long and Magerko, 2020) and digital skills (van Laar et al., 2017). One such novel job is the prompt engineer.
The emergence of LLMs and ‘prompt engineering’ (i.e., the practice of optimizing input prompts for LLMs) has sparked interest among researchers investigating how prompt engineering is applied in the job market, and the qualifications required for this emerging job role.
Kutela et al. (2023) examined the nature, qualifications, and compensation of prompt engineers through an analysis of LinkedIn job descriptions. Their study mainly focused on collocated keywords and frequency counts using statistical approaches. While these methods are well established and documented, they may fail to capture higher-level semantic meaning in the text and struggle to answer more complex questions about the profession’s requirements and expectations.
George et al. (2023) used a qualitative method of interviewing industry experts, university researchers, and students to examine opportunities for prompt engineers, including job requirements and the characteristics of these emerging roles. Although this approach is appropriate for the research questions, they provide no transparency regarding their interview protocols, participant selection criteria, or analytical framework. This lack of transparency throughout the research process raises questions about the credibility of their results.
With the rise of LLMs, significantly more researchers have approved and applied these models to qualitative research (Morgan, 2023; Kirsten et al., 2025; Zhang et al., 2024). Fecher et al. (2025) empirically explored the implications of large language models on the science system, including the required competencies for their effective use. Several studies (e.g., Kirsten et al. (2025); Zhang et al. (2024)) provide their own frameworks and approaches to effectively apply LLMs to qualitative research methods. This paper takes a mix of some of the suggested approaches and adapts them to our specific needs.
There is consensus among academic researchers, online communities, and industry professionals on the perishable nature of prompt engineering skills. Muktadir (2023) researched the historical development of prompting, which shows the rapid evolution of prompt engineering and the need for continuous adaptation of skills in this domain. In the following section, we review approaches for extracting skills from job postings.
2.2 Skill Extraction Techniques
The analysis of prompt engineering skills requires robust methodologies to identify the skills from job descriptions, a task that involves information extraction (IE) and natural language processing (NLP) techniques. This analysis faces multiple challenges, stemming from the lack of high-quality data about the nature of work (Frank et al., 2019), the unstructured nature of job descriptions (Ostrogonac et al., 2021), the inherent limitation of information retrieval and extraction techniques (Adnan and Akbar, 2019; Manning et al., 2008), and the evolving terminology within the prompt engineering field (Schulhoff et al., 2025).
Kivimäki et al. (2013) suggest a method called Elisit that uses classical information retrieval techniques. The method consists of two phases. First, similarity matching is used to map input text to Wikipedia articles using TF-IDF or other similarity measures. The method then applies a modified spreading activation algorithm on Wikipedia’s hyperlink graph to identify skill nodes. This approach would work with more established fields, because it relies on Wikipedia to have sufficient information. However, this approach may not be as effective for analyzing prompt engineering, an emergent field that is still largely unexplored and potentially underrepresented in established knowledge bases, such as Wikipedia.
Sateli et al. (2017) presented ScholarLens, a system for automatically constructing user profiles from scholarly literature, including competences. Their approach applies techniques from natural language processing (NLP) and linked open data entity linking.
Zhang et al. (2022a) presented Kompetencer, a Danish dataset of job postings with nested spans of competences, resulting from classifying job competences from job postings. The authors created this dataset by using the API of the European Skills, Competences, Qualifications and Occupations (ESCO; le Vrang et al. (2014)) taxonomy to obtain fine-grained labels via distant supervision. However, this taxonomy does not captured the novel set of skills required in the emerging role of prompt engineering.
Calanca et al. (2019) analyzed soft skills in job advertisements by drawing on a semi-automatic approach based on crowdsourcing and text mining for extracting a list of soft skills. Our work differs in that we do not employ crowdsourcing and also extract hard skills.
Michalczyk et al. (2021) specifically investigated the data scientist job role, by processing 25,104 online job postings with a text mining approach combining topic modeling, clustering, and expert assessment. The authors identified six job roles in data science in three major knowledge domains.
Figure 1: Skill extraction results using SkillSpan (Zhang et al., 2022b). The highlighted text spans show extracted skills with contextual information that complicates downstream analysis.
Zhang et al. (2022b) applied an approach based on sequence labeling. Their model, SkillSpan, is a BERT-based model that is fine-tuned on manually annotated job postings. Their approach reached 57.7% F1 score, and their approach also includes transparent annotation guidelines (Zhang et al., 2022b) and was later evaluated by other researchers (e.g., Nguyen et al. (2024)). However, their approach faces a critical limitation: because of the way they annotated skills, the extracted skills often include long spans containing extraneous information. As shown in Figure 1, the identified skills—while being relevant—include contextual elements complicating downstream analysis such as skill categorization.
To overcome the limitations in the above approaches, many researchers have turned to pre-trained language models, as reviewed in the following section.
2.3 Skill Extraction with LLMs
Recent research has shifted toward leveraging Large Language Models (LLMs) for skill extraction (Nguyen et al., 2024; Herandi et al., 2024), offering promising alternatives to traditional sequence labeling approaches. Notable among these approaches is the work by Herandi et al. (2024), who fine-tuned an 8B-parameter LLaMA model (Touvron et al., 2023) on the SkillSpan dataset (Zhang et al., 2022b), achieving a 64.8% F1 score using their specialized prompting techniques.
Gnehm et al. (2022) presented an approach for extracting and classifying skill requirements in German-speaking job postings. Their method adopts pre-trained transformer-based language models for extracting skills, using context from job postings and the ESCO taxonomy le Vrang et al. (2014). As with Kompetencer (Zhang et al., 2022a), the approach is limited as it draws on an existing domain taxonomy to match potentially novel skills.
Large language models have also peaked interest in research into leveraging LLMs in qualitative research. Morgan (2023) has examined the effectiveness of using ChatGPT in qualitative analysis. While the language model is less successful at locating subtle themes, it is more successful at reproducing concrete, descriptive themes. Kirsten et al. (2025) conducted interviews with 15 HCI researchers experienced in qualitative data analysis to understand their concerns about leveraging LLMs in research. Based on the research, they suggest frameworks to leverage the strengths of LLMs while minimizing problems that come with LLM use. Zhang et al. (2024) further advanced this field by developing a comprehensive framework for designing prompts to enhance ChatGPT’s effectiveness in thematic analysis. Through semi-structured interviews with seventeen participants and collaboration with thirteen qualitative researchers, they identified that improving transparency, providing guidance on prompts, and strengthening users’ understanding of LLMs’ capabilities significantly enhanced researchers’ ability to effectively utilize ChatGPT for qualitative analysis.
This paper adopts an LLM-based approach due to its demonstrated effectiveness, while avoiding sequence labeling methods to extract skills. By combining manual annotation with LLM-assisted qualitative analysis, as suggested by Kirsten et al. (2025), and incorporating prompt design principles from Zhang et al. (2024), we aim to overcome the limitations of traditional skill extraction techniques in analyzing the emerging field of prompt engineering. In the following section, we describe our approach to extracting skills from job postings.
3 Method
3.1 Data Collection
For this study, we use a dataset of job postings scraped from LinkedIn on April 7, 2025. The one-day crawl provides us a snapshot view into the online job market on this given date. While job postings on LinkedIn do not reflect all jobs advertised online, LinkedIn represents one of the largest global platforms for professional networking and job advertisements, making it a valuable source for analyzing current labor market trends. The data was crawled across the 60 countries with the most technological expertise, based on the U.S. News & World Report rankings (U.S. News, 2025). The countries are listed in Appendix A. The distribution of jobs reflects the concentration of tech jobs world-wide. For our comparative analysis, we selected four job roles for being highly relevant and potentially overlapping with the role of prompt engineer. These job roles are data engineer, data analyst, machine learning engineer, and data scientist.
The data collection was conducted in a 2 steps process: First we gathered all links to job postings that match the search criteria, then we extracted the detailed information for each job posting. For identifying job roles, we searched for exact matches in job titles, using quoted keywords to ensure precise matching jobs (e.g., “prompt engineer”). This approach allowed us to create a precise snapshot of job market for these five roles at a specific point in time.
3.2 Data Preprocessing
Prior to analysis, the dataset underwent two sequential preprocessing steps: first, we removed duplicate job postings to ensure each listing was unique (Section 3.2.1), and then we organized the remaining entries into coherent groups based on similarity (Section 3.2.2).
3.2.1 Deduplication
Some of the job postings in the data are duplicates or near-duplicates. For instance, international companies may create the same job postings across multiple regions. These duplicate job postings are sometimes not an exact match, but differ slightly in their content. This makes the deduplication process non-trivial.
To address this, we calculate the similarities between job postings and determine a cut-off value in similarity to filter out duplicate job postings. To this end, we first grouped the dataset by job roles (“data engineer”, “prompt engineer”, etc.) to reduce the computational complexity and improve accuracy. Then, for each group, we applied a three step process, consisting of vocabulary construction, TF-IDF calculation for each document, and pairwise cosine-similarity comparison.
  1. 1.
    Vocabulary construction: We first tokenized all documents and built a vocabulary (i.e., a set of all unique words across the entire corpus). The size of the vocabulary $v$ determined the dimensionality of the TF-IDF vectors.
  2. 2.
    TF-IDF vector for each document: Each job posting was represented as a sparse vector of length $v$. Each component in the vector corresponds to a word in the vocabulary. If a word did not appear in the document, its TF-IDF value is zero.
  3. 3.
    Cosine similarity: Given the vector representation of documents, we computed the cosine similarity between any two job postings by comparing their TF-IDF vectors.
Figure 2 displays the distribution of similarity scores across all job posting pairs. Based on visual inspection of this distribution, we can identify a distinct secondary mode at a threshold of 0.95 (highlighted with a dashed red line), above which job postings were considered duplicates.
Figure 2: Distribution of similarity scores between job posting pairs. The red vertical line indicates our chosen threshold (0.95) for identifying duplicate postings.
3.2.2 Grouping
Using connected component clustering, we identified groups of similar job postings and retained only one representative from each cluster. Connected component clustering is a graph-based method that groups items into clusters by linking pairs with similarity above a defined threshold, such that each cluster corresponds to a connected component in the resulting similarity graph. We created an undirected graph where nodes represent job postings, and edges connect nodes with similarity threshold above 0.95. We then identified all connected components using Depth-First Search (DFS) (Hopcroft and Tarjan, 1973). A connected component is a subgraph where there exists a path between any pair of nodes, meaning all job postings in the component are transitively similar to each other. From each connected component, we randomly selected a representative job posting for further analysis. This process reduced our initial dataset from 34,010 raw records to 20,644 unique job postings ($\approx 33\%$ reduction), including 72 prompt engineer positions. The following section describes how we extracted skills from these job postings.
3.3 Skill Extraction
To address RQ2 (What skills and qualifications are required for the role?), we initially attempted to apply the SkillSpan model (Zhang et al., 2022b) for skill extraction. However, this approach proved suboptimal for our specific dataset. The model frequently extracted entire phrases as skill entities (see Figure 1), making it hard to identify the skill in our analysis. Combined with the relatively low accuracy of the model when applied to our dataset, SkillSpan was not a feasible option to extract the skills from the job postings.
# Step 1: Examine the job description
You are an expert skill extractor with specialization in analyzing technical job postings. I’m going to provide you with a job description text. Your task is to carefully analyze this text to extract structured skill information.
# Context Specification
The input is a raw job posting that may contain various sections including job overview, responsibilities, requirements, and company information.
# Step 2: Extract and categorize information
Follow this step-by-step process:
1. Read the entire job description carefully
2. Identify all mentioned skills, qualifications, and requirements
3. Categorize each identified item into the appropriate category:
-- soft_skills: Interpersonal and behavioral attributes (e.g., communication, teamwork)
-- hard_skills: Technical abilities and specific knowledge (e.g., Python, prompt engineering)
-- qualifications: Educational background, certifications, or experience requirements (e.g., Bachelor’s degree, 3+ years experience)
-- other: Any requirements that don’t clearly fit the above categories
4. Determine the application domain where prompt engineering will be applied. This is NOT just ‘‘tech’’ or ‘‘AI’’ but the specific industry or field where the company will apply prompt engineering (e.g., healthcare, finance, education, marketing, software development)
# Output Format
Respond ONLY with a valid JSON object using this exact structure:
‘‘soft_skills’’: [‘‘skill1’’, ‘‘skill2’’, ...],
‘‘hard_skills’’: [‘‘skill1’’, ‘‘skill2’’, ...],
‘‘qualifications’’: [‘‘qual1’’, ‘‘qual2’’, ...],
‘‘job_domain’’: ‘‘The specific industry or application area where prompt engineering will be applied’’,
‘‘other’’: [‘‘item1’’, ‘‘item2’’, ...]
Important guidelines:
-- Extract skills as concise noun phrases (1--4 words typically)
-- Separate distinct skills (don’t combine multiple skills into one entry)
-- Format consistently using lowercase
-- Include ALL relevant skills mentioned, beyond prompting
-- For job_domain, identify the SPECIFIC industry or application area where the company operates and will apply prompt engineering (NOT just ‘‘technology’’ or ‘‘AI’’)
-- If multiple domains are mentioned, list the primary domain or focus
-- The output must be valid JSON (double quotes, commas between items)
-- Contain only the JSON object, no additional text
Figure 3: Prompt for skill extraction from job postings
Instead, we followed the qualitative analysis framework proposed by Zhang et al. (2024), which suggests specific techniques for designing effective LLM prompts for qualitative analysis of data. Figure 3 showcases our prompt, which has the following key elements:
  • Background understanding: We tell the LLM it was an “expert skill extractor” and explain what job postings look like. This helps the LLM better understand what it is analyzing.
  • Clear goal: We clearly state what we want – to identify and sort skills from job postings into categories, such as soft skills, hard skills, and qualifications. Being specific improved the results.
  • Step-by-step process: We give the LLM a clear process to follow: read the whole job post, find skills, and sort them into categories. This guided approach led to more consistent results.
  • Input description: We explained the input are job postings
  • Structured output: We asked for results in a specific format (JSON). The expectations of the output are described in detail to avoid any misinterpretation. This makes it easier to collect and analyze the data consistently across all job postings.
  • Expert role: By asking the LLM to act as an expert skill extractor. This has been found to be more effective (Dong et al., 2024)
  • Clean data: The data has been cleaned of unnecessary characters and change to lower case to ensure data uniform.
The extracted skills contained some variation in terminology and phrasing. For instance, “Python programming,” “Python development,” and “coding in Python” all refer to the same skill. To standardize the extracted information, we implemented an approach using k-means clustering to group semantically similar skills together. Each extracted skill was first mapped to a high-dimensional vector using sentence transformers (Reimers and Gurevych, 2019), specifically the “all-MiniLM-L6-v2” variant which builds upon the MiniLM architecture (Wang et al., 2020). This model creates embeddings that capture the semantics of text phrases, meaning that skills with similar meanings are located close to each other in the embedding space. We then applied k-means clustering to these skill vectors. K-means partitions the vector space into k clusters by minimizing the within-cluster variance, assigning each skill to the nearest cluster centroid. To determine an appropriate number of clusters, we conducted an iterative evaluation: we varied k from 5 to 20 and manually examined the top terms in each cluster to assess semantic coherence. Based on this evaluation, we selected 8 clusters for soft skills and 12 clusters for hard skills, as these provided the most interpretable groupings.
We visualize the embedding clusters with UMAP (McInnes et al., 2018), a dimensionality reduction technique that maps high-dimensional data (in our case, embeddings) to a two-dimensional space while preserving both local and global structures. UMAP was found to have advantages over techniques such as t-SNE and Principal Component Analysis (Coenen and Pearce, nd; McInnes et al., 2018).
This approach helped to draw a clearer picture of requirements of the different job roles. To name the cluster, we retrieved the top 5 skills (by term frequency) in each cluster and manually labeled each cluster based on these skills.
3.4 Correlation of Soft Skills and Hard Skills
In order to understand the relationships between soft skills and hard skills, we conducted an adjusted Pearson residual analysis using contingency tables (Agresti, 2013). This statistical approach allowed us to identify which combinations of soft skills and hard skills appear together more or less frequently than would be expected. To this end, we constructed contingency tables with hard skills as columns and soft skills as rows, where each cell values is adjusted Pearson residual. The adjusted Pearson residual for cell $(i,j)$ is calculated as:
$r_{ij}=\frac{O_{ij}-E_{ij}}{\sqrt{E_{ij}(1-p_{i+})(1-p_{+j})}}$ (1)
where $O_{ij}$ is the observed frequency in cell $(i,j)$, $E_{ij}$ is the expected frequency under independence, $p_{i+}$ is the marginal proportion for row $i$, and $p_{+j}$ is the marginal proportion for column $j$.
Adjusted Pearson residuals follow a standard normal distribution under the null hypothesis of independence. Values with $|r_{ij}|>2.0$ ($p<0.05$) indicate significant deviations from independence, with positive values indicating skills that appear together more often than expected and negative values indicating skills that appear together less often than expected.
4 Results
4.1 Descriptive Data Overview (RQ1)
After removing duplicates, our final dataset had 20,662 unique job postings across five job roles (prompt engineer, machine learning engineer, data scientist, data engineer, and data analyst). Figure 4 shows how these jobs were distributed in the dataset. Despite the high relevance of LLMs in the tech industry today, prompt engineer positions make up a tiny part of the job market (72 jobs), compared to other data and AI roles, such as data engineer (8,272 jobs), data analyst (6,782 jobs), data scientist (5,524 jobs), and machine learning engineer (1,958 jobs).
Figure 4: Distribution of job postings across roles after removal of duplicates
(a) Soft skills
(b) Hard skills
Figure 5: Skills cluster distribution for all analyzed jobs
This distribution confirms that while LLM use has grown quickly in industry, dedicated prompt engineer positions are still rare in today’s job market, making up less than 0.5% of related roles in our dataset. This matches our expectation that the role is still emerging and has not, yet, become as established as older positions.
4.2 Skills Analysis Across Job Roles
To understand what makes prompt engineering unique as a novel profession, we analyzed and compared both soft skills and hard skills required across different job roles. This analysis is contrasted with an analysis of the distinct skill profiles for prompt engineers (see Section 4.3), giving insights into how these roles differ despite working in related areas.
4.2.1 Soft skills requirements across jobs
Figure 5(a) and Table 2 show the results of our soft skill analysis across all collected jobs. Our k-means clustering identified eight distinct groups of soft skills. Statistical testing confirms significant differences between the skills profiles of the different job roles ($\chi^{2}(28,N=20662)=1377.78$, $p<0.0001$), validating that the skill requirements variation across roles is not due to random chance.
The soft skill distribution across all data roles shows that three clusters dominate the requirements: Adaptability & Critical Thinking (20.7%), Teamwork & Collaboration (19.7%), and Communication Skills (19.6%). Together, these three groups represent about 60% of all soft skills mentioned in job postings, showing a strong industry-wide focus on these interpersonal and thinking abilities.
Problem-Solving Abilities (12.5%) and Leadership & Independence (11.8%) form the second tier of important soft skills across data roles. The remaining clusters—Attention to Detail (5.8%), Stakeholder Management & Business Acumen (5.3%), and Time & Project Management (4.7%)—though mentioned less often overall, show important differences between specific job types, as we’ll examine in the following sections.
4.2.2 Hard skills requirements across jobs
7(b) and Table 4 depict the results of our comparative hard skill analysis across job postings. We identified twelve distinct hard skill clusters that appear consistently across job postings. Chi-square testing revealed highly significant differences among job types ($\chi^{2}(44,N=20662)=67291.55$, $p<0.0001$), confirming that the technical skill profiles across different job roles represent meaningful distinctions rather than random variation. This significant variation reflects the specialized technical requirements of each job position. Big Data Processing (25.1%) emerges as the most frequently required technical skill cluster across all job roles, followed by Data Warehousing & Modeling (12.6%) and SQL & Database Systems (8.7%). These three clusters form the foundation of technical requirements for data professionals, accounting for approximately 46% of all hard skills mentioned in job postings.
4.3 Skills Analysis for Prompt Engineers
4.3.1 Soft skills requirements for prompt engineers (RQ2)
Table 1: Soft skills clusters for Prompt Engineers
Cluster Top Skills (count) Freq.
Teamwork & Collaboration collaboration (27), teamwork (10), team collaboration (5), collaboration skills (5), team player (3) 21.9%
Communication Skills communication (15), communication skills (10), excellent communication skills (4), strong written communication (4), excellent communication (3) 21.9%
Adaptability & Critical Thinking creativity (7), adaptability (6), self-driven (2), creative mindset (2), innovation (2) 20.3%
Problem-Solving Abilities problem-solving (14), problem-solving skills (13), problem-solving mindset (3), analytical mindset (3), creative problem-solving (3) 15.8%
Leadership & Independence analytical skills (16), strong analytical skills (3), project management skills (2), organizational skills (2), ability to explain complex concepts (2) 10.3%
Stakeholder Management & Business Acumen ownership (2), business judgement (1), consultative culture (1), client collaboration (1), ownership and accountability (1) 4.2%
Attention to Detail attention to detail (7), documentation (1), exceptional attention to detail (1), quick learner (1), ability to identify patterns (1) 3.5%
Time & Project Management project management (3), excellent time management (1), ability to prioritize workload (1), fast execution mindset (1) 1.9%
Table 2: Soft skills clusters across other tech roles
Cluster Top Skills (count) Freq.
Adaptability & Critical Thinking adaptability (1509), creativity (757), curiosity (722), critical thinking (666), self-motivated (476) 20.7%
Teamwork & Collaboration collaboration (7462), teamwork (4225), team collaboration (1066), team player (796), collaborative (444) 19.7%
Communication Skills communication (8114), excellent communication skills (1163), communication skills (987), strong communication skills (986), interpersonal skills (961) 19.6%
Problem-Solving Abilities problem-solving (5053), problem-solving skills (1117), problem solving (974), analytical thinking (775), strong problem-solving skills (477) 12.5%
Leadership & Independence analytical skills (1672), mentoring (1068), leadership (1042), ability to work independently (782), mentorship (665) 11.8%
Attention to Detail attention to detail (1883), detail-oriented (450), strong attention to detail (148), results-oriented (89), detail oriented (72) 5.8%
Stakeholder Management & Business Acumen stakeholder management (459), ownership (365), relationship building (247), stakeholder engagement (229), business acumen (165) 5.3%
Time & Project Management time management (397), project management (391), prioritization (99), multi-tasking (95), organized (87) 4.7%
(a) Soft skills
(b) Hard skills
Figure 6: Comparative distribution of skills clusters across job roles
As depicted in 6(a), 7(a) and Table 2, prompt engineer roles place a particularly strong emphasis on collaborative communication, which together accounts for 43.8% of their soft-skill requirements—higher than the $\approx 39\%$ observed in other roles. This reflects how prompt engineers must work closely with stakeholders to translate practical needs into effective language-model prompts.
Beyond teaming and communication, creative problem-solving stands out: terms such as “creative problem-solving” and “analytical mindset” appear among the top skills for prompt engineers, underscoring the innovative thinking needed to design and refine prompts. The frequent mention of strong written communication further highlights the text-centric nature of the prompt engineer’s work.
By contrast, time and project management skills are cited far less often in prompt engineer postings (1.9% vs. 4.7% across other roles), suggesting that rapid prototyping and iterative experimentation take priority over formal scheduling. When analytical abilities are required, they are most often linked with an “ability to explain complex concepts,” pointing to the interpretive and educational facets of the role as prompt engineers translate model behavior into actionable insights.
4.3.2 Hard skills requirements for prompt engineers (RQ3)
When examining the hard skills specifically required for prompt engineers, a distinctive technical profile emerges that highlights the unique requirements of this emerging role. Figure 7 presents the hard skills distribution specific to prompt engineer positions. While data scientists and machine learning engineers focus on model development and statistical analysis, prompt engineers show the strongest emphasis on Machine Learning & AI (22.8%), specifically relating to large language models, natural language processing, and machine learning frameworks. Also notable is the high importance of Agile & Testing skills (18.7%), ranking second for prompt engineers but appearing much lower for other job roles (3.6% across all roles). This category for prompt engineers is focused on skills directly related to prompt engineering techniques (47 mentions), prompt design (6 mentions), and chain-of-thought prompting (4 mentions), highlighting the specialized nature of this role compared to traditional testing methods used by other professions. Big Data Processing remains important (17.8%), but with a specific focus on tools like PyTorch (10), Hugging Face (4), and LangChain (4), which are designed for working with language models rather than for traditional analytics. The importance of ETL & DevOps (10.8%) with emphasis on version control, Node.js, and API integration reflects the need for integrating LLM capabilities into production systems. Python & Scripting (9.2%) shows the continued importance of Python (33 mentions) as the main programming language for prompt engineers, similar to other job roles, though with more specific focus on developing code that interfaces with language models.
Table 3: Hard skills clusters for Prompt Engineers
Cluster Top Skills (count) %
Machine Learning & AI large language models (22), natural language processing (16), tensorflow (10), machine learning (9), machine learning frameworks (5) 22.8
Agile & Testing prompt engineering (47), prompt design (6), chain-of-thought prompting (4), a/b testing (3), writing prompts (3) 18.7
Big Data Processing pytorch (10), typescript (4), hugging face (4), langchain (4), nlp (4) 17.8
ETL & DevOps version control (4), node.js (4), api integration (3), chatbot development (3), automation tools (3) 10.8
Python & Scripting python (33), python programming (7), programming languages (4), python development (3), scripting (3) 9.2
Performance Optimization retrieval-augmented generation (4), statistical modeling (2), instruction tuning (2), scalability (2), optimizing prompts (2) 7.6
Cloud Infrastructure azure (5), aws (4), cloud platforms (4), microsoft azure (3), cloud-based application development (2) 4.9
Data Warehousing & Modeling data analysis (14), data science (2), data manipulation (2), dataset creation (1), data transformation (1) 4.1
SQL & Database Systems sql (4), vector databases (3), database management (2), vector db (1), postgresql (1) 2.0
Data Visualization & BI data visualization (6), data visualization tools (3), google analytics (1), analytics (1), visualization tools (1) 1.7
Distributed Computing big data tools (1), pyspark (1), databricks (1) 0.4
Data Pipeline Development rag pipelines (1) 0.1
Table 4: Hard skills clusters across other tech roles
Cluster Top Skills (count) %
Data Science & Analytics Tools r (2578), power bi (2495), spark (2488), tableau (2479), pytorch (1851) 25.1
Data Warehousing & Modeling data analysis (4279), data modeling (2945), data engineering (1690), data warehousing (1517), data governance (1012) 12.6
SQL & Database Systems sql (11247), nosql (644), postgresql (542), mysql (416), mongodb (371) 8.7
Performance Optimization statistical analysis (1156), data mining (865), predictive modeling (781), statistical modeling (750), computer vision (551) 8.2
Cloud Infrastructure aws (2577), azure (1473), kubernetes (1412), azure data factory (851), terraform (801) 7.6
Machine Learning & AI machine learning (4268), tensorflow (1820), deep learning (1166), natural language processing (760), generative AI (551) 7.4
Python & Scripting python (12616), python programming (409), matplotlib (248), programming in python (239), shell scripting (180) 7.1
ETL & DevOps etl processes (1341), devops (475), github (354), etl tools (339), c# (228) 6.9
Distributed Computing databricks (1884), pyspark (1663), hadoop (1355), apache spark (862), bigquery (698) 5.5
Data Visualization & BI data visualization (2983), excel (1198), data analytics (827), data visualization tools (701), predictive analytics (300) 5.3
Agile & Testing a/b testing (579), prompt engineering (285), agile methodologies (252), project management (130), agile methodology (124) 3.6
Data Pipeline Development data pipelines (955), ci/cd pipelines (521), data pipeline development (415), etl pipelines (392), data pipeline design (207) 1.9
(a) Soft skills
(b) Hard skills
Figure 7: Skills cluster distribution for prompt engineer positions, in comparison to skills in other jobs
Figure 6(b) and tables 4 and 4 present a side-by-side comparison of hard skill requirements across all five job roles, further illustrating the distinctive technical profile of prompt engineers. This comparison reveals several key insights:
  • Prompt engineers require the highest proportion of Machine Learning & AI skills (22.8%) compared to machine learning engineers (21.6%) and data scientists (16.0%), yet their focus is specifically on language models rather than general machine learning algorithms.
  • Prompt engineers show a significantly higher demand for Agile & Testing skills (18.7%) compared to all other roles, with data analysts being the next highest at only 8.5%. However, the specific skills within this cluster differ dramatically, with prompt engineers focusing on prompt design methods rather than traditional software testing approaches.
  • The emergence of specialized tools such as LangChain, Hugging Face, and vector databases in prompt engineer job descriptions highlights the development of a specialized toolset for this role that differs from traditional data engineering and machine learning tools.
  • Data pipeline development, which is critical for data engineers (3.1%), is almost nonexistent for prompt engineers (0.1%), suggesting less focus on traditional ETL workflows in the prompt engineering role and more emphasis on specialized prompt optimization processes.
The emergence of these specialized skill requirements confirms that prompt engineering is developing as a distinct technical discipline rather than simply being an extension of existing job roles. This technical specialization, combining AI expertise with prompt-specific methods and tools, supports our finding that prompt engineering represents a new professional category with its own distinct skill profile, even as it remains a relatively small segment of the current job market.
4.3.3 Correlation between soft and hard skills of prompt engineers
Figure 8 reveals nuanced patterns that help us understand what employers value in the role of prompt engineers beyond general skill listings. The significant correlation between Teamwork & Collaboration and ETL & DevOps ($r=2.22$) indicates that prompt engineers working on data integration and deployment pipelines are expected to collaborate extensively, likely because these tasks involve cross-functional workflows. There further is a positive association between the skills Leadership & Independence and Machine Learning (ML) & AI ($r=2.21$). This may suggest that high autonomy in ML-heavy prompt engineering roles is required, possibly due to the experimental and exploratory nature of prompt design for LLMs. Further, Attention to Detail is often paired with SQL & Database Systems skills ($r=2.51$). Data-related roles for prompt engineers require precision, likely for accurate data extraction or grounding prompts in structured databases. The significant co-occurrence of Time & Project Management and Data Warehousing & Modeling ($r=2.12$) could suggest that prompt engineers working in structured data environments also manage timelines or project deliverables. Engineers optimizing prompts for latency or token efficiency need to think critically and iterate quickly, which could explain the co-occurrence of the skills Adaptability & Critical Thinking and Performance Optimization ($r=1.89$).
We identified two significant negative co-occurrences of skills. The skills Teamwork & Collaboration are negatively correlated with SQL & Database Systems ($r=-2.23$). This may imply that prompt engineers working on back-end–focused and database-oriented tasks are less integrated with team processes. The second negative association is between Stakeholder Management & Business Acumen and Machine Learning & AI ($r=-2.53$). This could indicate that prompt engineers prioritize business-facing tasks over heavy technical AI work. It could, however, also indicate a potential technical sub-role within the prompt engineer role, where machine learning and developer-oriented tasks are emphasized over higher-level communication and business tasks.
Figure 8: Skills co-occurrence between soft and hard skills for prompt engineer positions. High values indicate pairs of skills that are often mentioned together in job descriptions, whereas low values indicate the opposite. Values of $|r|>2.0$ indicate significant deviations from expected values ($p<0.05$).
5 Discussion
In this section, we discuss our findings on the emergence of prompt engineering as a profession, its skill requirements, and implications for the job market.
5.1 The Emerging Status of Prompt Engineering
Our analysis shows that prompt engineering remains in its early stages as an emerging profession. With just 72 prompt engineer positions identified (less than 0.5% of the related roles in our dataset), the role has not yet achieved widespread adoption across industries. This limited presence confirms that despite significant industry attention on large language models, dedicated prompt engineering positions are still relatively rare. This finding aligns with the timeline of LLM development and adoption. While LLMs have gained rapid attention since late 2022, the establishment of specialized roles typically lags behind technology adoption as organizations first experiment with assigning these new responsibilities to existing staff before creating dedicated positions.
5.2 The Distinct Soft Skill Profile of Prompt Engineers
While skills may become obsolete (Horton and Tambe, 2025), new skills can emerge, giving rise to professions with a distinct skill profile. This skill profile has a strong overlap with what has been termed the “21st-century” skill profile in prior literature (van Laar et al., 2017; Federiakin et al., 2024). For instance, van Laar et al. (2017) identified seven 21st-century skills in their systematic literature review: technical, information management, communication, collaboration, creativity, critical thinking, and problem solving. This set of skills is well represented in our analysis of the prompt engineering profession. The two key soft skills identified for the emerging role of prompt engineers are an enhanced focus on communication (21.9%) and creative problem-solving (15.8%). Prompt engineer positions emphasize strong communication skills more than the other investigated job roles. This reflects the expectation that prompt engineers must effectively communicate their innovative approaches and solutions to various stakeholders. Prompt engineers are also expected to apply creative approaches to problem-solving, with skills such as “creative mindset” and “creative problem-solving” appearing frequently in job postings. This suggests employers recognize that prompting is an innovative field that requires engineers to think creatively when solving problems.
Our analysis reveals that while communication and teamwork capabilities are universal soft skills in tech roles—aligning with Lievens and Sackett (2012)’s findings on interpersonal skills—prompt engineers show distinctively high proportions of these requirements. As 6(a) illustrates, Communication Skills and Teamwork & Collaboration each represent 21.9% of soft skill requirements, with Problem-solving Abilities (15.8%) also emphasized more strongly compared to other job roles. Interestingly, Time & Project Management skills appear least frequently (1.9%), which may align with prompt engineering being an innovative emerging field that lacks standardized management processes. These findings indicate that prompt engineers serve as innovation communicators at the interface between LLM capabilities and stakeholder needs.
Our analysis also shows that the skills profile of prompt engineers is only partially reflected in existing skills taxonomies, such as the UK skills taxonomy (Djumalieva and Sleeman, 2018) and the Burning Glass skills taxonomy (Burning Glass Technologies, 2019). Prompt engineers have a specific profile that not only leverages advanced linguistic expertise—including nuanced control of semantics, syntax, and discourse strategies—to iteratively craft and optimize prompts that effectively steer AI behavior, a dimension largely absent from traditional skills frameworks. In addition, prompt engineers, use tools and frameworks with scripting languages, as demonstrated by our analysis of hard skills. In the following section, we expand our discussion on the technical skills profile of prompt engineers.
5.3 Technical Skills Defining Prompt Engineering
The analysis of hard skills provides strong evidence that prompt engineering is emerging as a distinct technical discipline rather than simply an extension of existing job roles. Chi-square testing revealed highly significant differences among the skill profiles of different job types ($\chi^{2}(44,N=20662)=67291.55$, $p<0.0001$), confirming that these technical skill profiles represent meaningful distinctions. Notably, the side-by-side comparison with other job roles (cf. Table 4, Table 4, and 7(b)) highlights that prompt engineers stand out in four technical categories:
  1. 1.
    Machine Learning & AI (22.8%): According to our analysis, prompt engineers need broader AI skills than machine learning engineers (21.6%). Job postings often mention large language models (22 times), natural language processing (16), and AI tools for language work.
  2. 2.
    Agile & Testing (18.7%): Prompt engineers need significant more testing skills than other roles. Data analysts, for instance, only require this skill in 8.5% of the job postings. This matches the findings by Liu and Chilton (2022) – the prompt engineering process involves extensive testing and optimization.
  3. 3.
    ETL & DevOps (10.8%): Skills such as version control and API integration are important. This shows companies want to connect LLMs to their existing systems and workflows. This aligns with our experience when manually labeling a small subset of the job postings in our initial exploration of the data.
  4. 4.
    Emerging Tools and Domain-specific Methods (18.7%): Job postings often mention prompt engineering techniques (47 times), prompt design (6), and chain-of-thought prompting (4). These are special methods specific to this job role. Requirements for specific tools, such as LangChain (4), Hugging Face (4), and vector databases (3), show a unique toolkit is forming for prompt engineers.
In addition, prompt engineers require Python & Scripting (9.2%) skills. While Python is often used for ETL, statistical analysis, machine learning pipelines, or data visualization in traditional job roles, prompt engineers use Python specifically to interface with and manipulate large language models (LLMs). For instance, prompt engineers may use Python to orchestrate API calls to LLMs, construct dynamic prompts, parse model outputs, and evaluate generations – tasks that are more about language manipulation and real-time interaction than batch data processing. Prompt engineers also write lightweight scripts to prototype interactions, automate prompt variation, or integrate LLMs into applications, often with Python-based tools and libraries, such as requests, langchain, or openai.
Interestingly, Cloud Skills (4.9%) were found to be less important than for other related roles. This may be because prompt engineering is often positioned closer to application-level prototyping and model interaction rather than infrastructure deployment or scalable production workflows. Prompt engineers frequently work with hosted APIs and specific online tools (e.g., OpenAI, Anthropic, Cohere) rather than deploying models themselves, while cloud infrastructure (e.g., AWS, GCP, Azure) is more relevant for backend developers, machine learning engineers, or MLOps roles responsible for running and scaling models, data pipelines, or inference services. As a result, prompt engineers often operate in environments abstracted from cloud infrastructure, relying on tools that hide those complexities (e.g., notebooks, playgrounds, or SDKs). Thus, while cloud proficiency can be useful, it is typically not a central requirement in prompt engineering job descriptions.
This unique mix of skills shows prompt engineering is becoming a specialized field focused on making language models work well in real applications.
5.4 Implications for the Job Market
The emergence of prompt engineering as a distinct profession with specialized skill requirements has several implications for both job seekers and employers:
  1. 1.
    Hybrid skill set requirements: Prompt engineers need a unique combination of technical AI knowledge, strong communication, and strong creative skills. This hybrid profile is rare, as it blends competencies from traditionally separate domains—such as software engineering, linguistics, and design. In practice, prompt engineers must be capable of understanding model behavior, crafting precise language instructions, and iteratively refining outputs based on feedback. For employers, this means recruiting for a complex and still evolving role that few candidates are formally trained for, making talent acquisition and role definition more difficult.
  2. 2.
    Professional development pathways: The distinct skill profile of prompt engineering creates new opportunities for professional specialization, especially for individuals already working in adjacent fields. Data scientists, NLP practitioners, UX designers, and machine learning engineers may find in prompt engineering a logical progression or pivot point, particularly as organizations expand their use of large language models. The role offers a way to deepen one’s focus on LLM applications while also contributing to product design, user experience, or applied AI research. As demand grows, we may see the emergence of formal career tracks and professional certifications centered on prompt engineering.
  3. 3.
    Educational gaps: Existing academic programs in computer science and data science generally do not cover the core competencies required for prompt engineering. While students may gain exposure to machine learning and natural language processing, they typically receive little to no training in prompt design, model behavior analysis, or the human-in-the-loop techniques often used in real-world deployments. As a result, graduates entering the workforce may lack formal training on LLM-specific tools, evaluation practices, and interaction patterns. Addressing this gap will likely require curriculum innovation, including interdisciplinary courses that bridge AI, language, and design.
5.5 Limitations and Future Work
While our study provides valuable insights into the prompt engineering profession, we acknowledge several limitations:
  1. 1.
    Sample size: With only 72 prompt engineer positions identified on LinkedIn, our analysis gives insights into the emerging landscape of prompt engineering in an important online job market, but the result may not be indicative of wider developments.
  2. 2.
    Search keywords: We specifically searched for “prompt engineer” as a job role. However, some job advertisements may mention prompt engineering as a skill. These advertisements are not captured by our analysis, because our key subject of study is the role of prompt engineer as a separate job role.
  3. 3.
    Data processing: Specifically the clustering process this may still contains error because the clustering process are not perfect as the embedding is missing the context. We need an embedding that’s more context aware or a different clustering/grouping method.
  4. 4.
    Skills extraction: In our work, we focused on extracting soft and hard skills. This analysis could be expanded into extracting competencies, which include skills, knowledge, attitudes, and abilities (World Economic Forum, 2021). Such an analysis would provide a more nuanced understanding of the job profile of the prompt engineer.
  5. 5.
    Geographic limitations: Our data collection focused on LinkedIn job postings in countries with technological expertise, potentially missing regional variations in how prompt engineering roles are defined.
  6. 6.
    Temporal dynamics: Given the rapid evolution of LLM technology and methods, the skill requirements for prompt engineers are likely to change quickly, requiring regular update.
Future research could extend the analysis to a broader dataset, and also track how prompt engineering job descriptions evolve over time as the field matures. Further, as mentioned in Section 4.3.3, there could be distinct prompt engineering sub-roles focused more on technical or business aspects. Future work could investigate this further. Additionally, qualitative studies involving interviews with prompt engineers and their employers would provide deeper insights into the day-to-day responsibilities and challenges of this role. Finally, expanding the analysis to include salary data would help assess how the market values these specialized skills compared to traditional job roles.
6 Conclusion
Our data-driven analysis of prompt engineering job postings reveals the role of prompt engineer remains rare, comprising less than 0.5% of related positions despite growing LLM adoption. The emerging profession shows a distinctive skill profile combining strong communication and creative problem-solving with specialized AI knowledge. Most surprisingly, prompt engineering extends far beyond crafting textual prompts. The emphasis on testing and DevOps skills indicates that prompt engineers are professionals responsible for the entire solution lifecycle—designing, testing, integrating, and deploying LLM systems into production workflows. This creates both opportunities and challenges, with important caveats about the job role’s future. The combination of slow organizational adoption and rapid AI evolution creates uncertainty about the longevity of dedicated prompt engineering positions. As LLMs improve and no-code interfaces advance, some current prompt engineering tasks may become automated. Job seekers should develop broader AI foundations alongside prompt engineering skills, while employers may favor upskilling existing employees over creating specialized roles.
As LLM technology matures, we expect prompt engineering to evolve significantly. Future research should track changes in job descriptions, skills, and compensation to better understand how organizations integrate AI technologies and adapt to technological innovation.
Appendix
Appendix A List of Countries
United States (6553; 31.7%), India (3747; 18.1%), United Kingdom (1490; 7.2%), Canada (621; 3%), Singapore (514; 2.5%), Poland (494; 2.4%), Portugal (440; 2.1%), Spain (439; 2.1%), Germany (434; 2.1%), Philippines (418; 2%), Australia (347; 1.7%), Mexico (276; 1.3%), Israel (273; 1.3%), Netherlands (269; 1.3%), Romania (210; 1%), Brazil (209; 1%), France (199; 1%), Malaysia (195; 0.9%), Belgium (183; 0.9%), Ireland (176; 0.9%), Indonesia (160; 0.8%), China (152; 0.7%), Vietnam (141; 0.7%), Sweden (140; 0.7%), Colombia (136; 0.7%), South Africa (134; 0.6%), Italy (127; 0.6%), Hungary (122; 0.6%), Argentina (122; 0.6%), Egypt (117; 0.6%), Greece (114; 0.6%), Japan (109; 0.5%), Thailand (102; 0.5%), Ukraine (98; 0.5%), Czech Republic (97; 0.5%), Turkey (96; 0.5%), United Arab Emirates (93; 0.5%), Lithuania (89; 0.4%), Bulgaria (88; 0.4%), Switzerland (88; 0.4%), Taiwan (87; 0.4%), Pakistan (72; 0.3%), Saudi Arabia (70; 0.3%), Denmark (58; 0.3%), Finland (49; 0.2%), Serbia (47; 0.2%), Cyprus (45; 0.2%), Slovakia (39; 0.2%), New Zealand (38; 0.2%), Morocco (34; 0.2%), Nigeria (29; 0.1%), Luxembourg (29; 0.1%), South Korea (27; 0.1%), Sri Lanka (27; 0.1%), Estonia (26; 0.1%), Austria (26; 0.1%), Malta (22; 0.1%), Norway (17; 0.1%), Kenya (17; 0.1%), Croatia (17; 0.1%), Latvia (14; 0.1%), Bangladesh (14; 0.1%), Chile (13; 0.1%), Slovenia (11; 0.1%), Qatar (11; 0.1%), Tunisia (6; $<$0.1%), Russia (2; $<$0.1%), Belarus (2; $<$0.1%), Liechtenstein (1; $<$0.1%)